# Data-Enabled Predictive Control and Guidance for Autonomous Underwater Vehicles

### v1

@@ 5 Data-Enabled Predictive Control
The DeePC framework, introduced in (Coulson et al., 2019), relies on measured input–output trajectories that are sufficiently informative to capture the system dynamics, a property known as persistency of excitation (Willems et al., 2005). According to Willems’ Fundamental Lemma, any trajectory of a controllable and observable LTI system can be expressed as a linear combination of previously recorded input–output data. The collected input and output sequences, $u_{\mathrm{data}}$ and $y_{\mathrm{data}}$, are arranged into block Hankel matrices $\mathscr{H}(u)$ and $\mathscr{H}(y)$ with $r$ block rows and $c$ columns, as defined in (Coulson et al., 2019). Alternatively, the data can be structured as a Page matrix $\mathscr{H}_{P}(u)$ as in  (21), which groups consecutive columns into separate pages (Coulson et al., 2021).
@@
General persistency of excitation requires these data matrices to be of full row rank (Coulson et al., 2019). For noisy data, a quantitative characterization of this condition was introduced in (Coulson et al., 2022), relating the degree of excitation to the singular value of the data matrix. For the DeePC formulation, the data matrices are partitioned into past and future components as in (22), where $(U_{p},Y_{p})$ encode past trajectories for initialization, and $(U_{f},Y_{f})$ define the future trajectories used for prediction.
@@
Here, $g\in\mathbb{R}^{T-T_{\mathrm{ini}}-N+1}$ is the decision vector that linearly combines past trajectories to reconstruct the system behavior, and $r\in\mathbb{R}^{p}$ denotes the reference trajectory. $Q\in\mathbb{R}^{p\times p}$ and $R\in\mathbb{R}^{m\times m}$ are positive definite weighting matrices. The regularization terms ${\|g\|_{\lambda_{g}}^{2}}$ and ${\|\sigma_{y}\|_{\lambda_{y}}^{2}}$, with $\lambda_{g},\lambda_{y}>0$, follow the approach of (Huang et al., 2023) to improve robustness against measurement noise and nonlinearities. Input and output constraints are directly incorporated through the box constraints and admissible sets $\mathcal{U}$ and $\mathcal{Y}$. Theoretical and experimental guarantees regarding robustness and stability are provided in (Berberich et al., 2020, 2021; Berberich and Allgöwer, 2024).
@@ 6.2 DeePC Depth Control
While the data collection process for the inner loop is straightforward as for the heading control, collecting suitable data for the outer loop requires additional consideration, as its effective plant consists of the inner DeePC controller combined with the AUV dynamics. At first glance, this appears incompatible with the DeePC framework, as the inner DeePC involves an embedded optimization, breaking the deterministic LTI structure typically assumed by Willems’ Fundamental Lemma. To overcome this, a perfect control assumption, analogous to the unity-gain property in cascaded PID control (Beard and McLain, 2012), is introduced. This idealized condition assumes that the inner loop perfectly tracks its reference, i.e., $\theta=\theta_{d}$, thereby allowing the pitch angle $\theta$ to be used as the effective input and the measured heave position $z_{p}$ as the output for the outer-loop data sequences.
@@ 8 Results
This section presents the simulation results obtained for the proposed DeePC-based control framework and compares its performance with the classical PI/PID control architecture across various scenarios, including nominal operation, external disturbances, nonlinear dynamic conditions and 3-D path following. For the classical architecture in all following scenarios, the best-performing PI/PID gains were selected from the preinitialized MSS toolbox for depth control and the tuning proposed in (Fossen, 2023) for heading control. The gain configuration of (Fossen and Aguiar, 2024) was also tested but yielded inferior performance. The PI/PID gains are as follows: $K_{p,\theta}=5.0,K_{i,\theta}=0.3,K_{d,\theta}=2.0,K_{p,\psi}=7.5,\\ K_{i,\psi}=0.75,K_{d,\psi}=15,K_{p,z}=0.1,K_{i,z}=1e-3$
@@ 8.1 General Control Performance
The PI/PID and DeePC-based controllers for heading and depth control are compared, with reference commands smoothed using the built-in LP filters of the MSS toolbox (Fossen and Perez, 2004). At $t=10\,\mathrm{s}$, a step command was applied toward a depth of $30\,\mathrm{m}$ and a heading of $60^{\circ}$. At $t=100\,\mathrm{s}$, a sinusoidal signal was superimposed on both reference trajectories to evaluate performance under more dynamic, albeit less realistic, reference variations. In both control approaches, the propeller speed was fixed to $n=1000\,\mathrm{rpm}$ during data collection and control execution. The specific DeePC hyperparameters (HP) are listed in Table 2.
@@ 8.3 Robustness Evaluation under Nonlinear Dynamics
8.5 Hyperparameter
The DeePC HPs were tuned in a two-stage process to ensure robust performance across the AUV operating range. In the first stage, broad parameter intervals were explored, followed by a refined search in a second stage, as detailed in Appendix E. The selected HPs, summarized in Table 2, were chosen based on the RMSE and the smoothness of the control inputs. For the 3-D path-following case in Section 8.4, the $\theta$-DeePC controller uses slightly modified weights, specifically $Q=100$ and $R=100$.
@@ 9 Conclusion and Further Work
This work successfully demonstrated DeePC for AUV control in both heading and depth regulation. A cascaded DeePC architecture was developed for depth control, incorporating a loop-frequency separation concept that proved effective under disturbance and robustness tests. Furthermore, the PALOS algorithm was proposed and integrated with DeePC, enabling successful 3-D waypoint path following of the REMUS 100 AUV. The results confirm that the proposed PALOS-DeePC approach constitutes a robust, data-driven, and predictive control solution, achieving high performance without requiring an explicit dynamic model and the associated effort. DeePC consistently outperformed the classical PI/PID control architecture, and although more aggressive gain tuning could narrow the performance gap, the predictive nature of DeePC inherently offers superior anticipatory behavior compared to reactive controllers.
Future work will extend the approach towards more nonlinear operation across the full speed range with gain-scheduled DeePC formulations, and advancing from waypoint-based path following to full trajectory tracking. Experimental validation on a real AUV would demonstrate the method’s practical applicability.
@@
Appendix A Additional Notation
To simplify the representation of the AUV dynamics, the state vector is expressed with respect to two coordinate systems: the North-East-Down (NED) frame and the body-fixed frame. Using the Euler angle transformation, the rotation matrix ${R}_{b}^{n}$ defined in (26) transforms velocities from the body-fixed frame $b$ to the NED frame $n$, as shown in (25(Fossen, 2021), where $v_{n}$ and $v_{b}$ are velocities with respect to the NED and body-fixed frames respectively. The inverse matrix ${R}_{n}^{b}=({R}_{b}^{n})^{-1}$ enables the transformation in the opposite direction. Here, $c\psi$ and $s\psi$ denote the cosine and sine of the angle $\psi$, respectively.
${v}_{n}={R}_{b}^{n}{v}_{b}$ (25)
$T_{k}=\begin{bmatrix}1&\sin\phi\tan\theta&\cos\phi\tan\theta\\[4.0pt] 0&\cos\phi&-\sin\phi\\[4.0pt] 0&\sin\phi/\cos\theta&\cos\phi/\cos\theta\end{bmatrix}$
Appendix B Parameter Introduction for AUV dynamic matrices
Prior to describing the components of the 6-DOF rigid-body equations of motion, some fundamental concepts regarding the vehicle geometry must be established. The REMUS 100 AUV is represented using the Myring hull equations (Myring, 1976) and, in the MSS toolbox, approximated as a spheroid (Fossen and Perez, 2004). The spheroid is defined by its longitudinal, transverse, and vertical semi-axes $a$, $b$, and $c$. Owing to the near-cylindrical shape of the vehicle, the transverse and vertical semi-axes are equal ($b=c$), and the geometry satisfies the relation in (30) in terms of the AUV’s length $L_{\mathrm{AUV}}$ and diameter $D_{\mathrm{AUV}}$. To ensure consistency with the vehicle mass, a rescaling of the semi-axes is performed according to (31), where $m$ is the vehicle mass and $\rho$ the fluid density. Solving these relations yields the semi-axis values $a$ and $b$ defining the equivalent spheroidal representation of the REMUS 100.
$\frac{L_{\mathrm{AUV}}^{2}}{a^{2}}+2\frac{D_{\mathrm{AUV}}^{2}}{b^{2}}=1$ (30)
$m=\frac{4}{3}\pi\rho ab^{2}$ (31)
$I_{x_{b}}=\frac{2}{5}m\,b^{2},$ (32)
$I_{y_{b}}=I_{z_{b}}=\frac{1}{5}m\,(a^{2}+b^{2}),$ (33)
$e=1-\left(\frac{b}{a}\right)^{2}$ (34)
$\beta_{0}=\frac{1}{e^{2}}-\frac{1-e^{2}}{2e^{3}}\ln\frac{1+e}{1-e}$ (36)
Following the $k$-factors introduced in (Lamb, 1924), the added-mass contributions to the system inertia can be estimated. The factors are defined as
yielding the diagonal added-mass matrix
While classical hydrodynamics predicts zero added moment in roll (element $(4,4)$ of $M_{A}$) for a spheroidal hull (Imlay, 1961), practical vehicles such as the REMUS 100 experience additional hydrodynamic contributions from the control surfaces and the propeller. Accordingly, this roll-related entry of the added-mass matrix is nonzero, $M_{A}(4,4)=r_{44}I_{x}$, capturing these effects in the simulation (Fossen and Perez, 2004).
$U_{r}=\sqrt{u_{r}^{2}+v_{r}^{2}+w_{r}^{2}}$ (39)
$U_{rh}=\sqrt{u_{r}^{2}+v_{r}^{2}},\quad U_{rv}=\sqrt{u_{r}^{2}+w_{r}^{2}}$ (40)
Appendix C The AUV dynamic matrices
In the MSS toolbox, $M_{A}$ is assumed constant, resulting in a constant $M$. The REMUS 100 exhibits both port–starboard and fore–aft symmetry, which simplifies the inertia representation according to Fossen (2021) to the sparse structured form shown in (44).
$C(\nu_{r})=C_{RB}(\nu_{r})+C_{A}(\nu_{r}).$ (45)
[…]
$D(\nu_{r})=-\operatorname{diag}\Big\{\,X_{u}e^{-3U_{r}},\;Y_{v}e^{-3U_{r}},\;Z_{w},\;K_{p},\;M_{q},\;N_{r}\,\Big\}.$
as well as lift forces in sway and heave, respectively,
$F_{\mathrm{lift}}=\tfrac{1}{2}\rho U_{r}^{2}\,S\,C_{L}(\alpha),\qquad F_{\mathrm{drag}}=\tfrac{1}{2}\rho U_{r}^{2}\,S\,C_{D}(\alpha),$ (55)
$\tau_{LD,Z}$ $=-F_{\mathrm{drag}}\sin\alpha\;-\;F_{\mathrm{lift}}\cos\alpha.$
$\tau_{LD}=\left[\tau_{LD,X},0,\tau_{LD,Z},0,0,0\right]^{\top}$ (57)
The cross-flow drag forces and moments are modeled following the strip-theory formulation by Faltinsen (1993) and Fossen (2021). The AUV hull is approximated as a series of two-dimensional strips of draft $T=D_{\mathrm{AUV}}$ and length $L=L_{\mathrm{AUV}}$, each experiencing a local flow velocity dependent on the sway and heave components $(v_{r},w_{r})$ and the corresponding angular rates $(r,q)$. The sectional drag is characterized by the two-dimensional drag coefficient $C_{\mathrm{d}}^{2\mathrm{D}}(v_{r},w_{r})$, yielding the cross-flow drag integrals in (58). The total generalized cross-flow drag vector acting on the vehicle is given in (59), representing the hydrodynamic forces and moments in sway, heave, pitch, and yaw due to lateral and vertical flow separation along the hull.
$\tau_{CF}=\left[0,\,Y_{CF},\,Z_{CF},\,0,\,M_{CF},\,N_{CF}\right]^{\top},$ (59)
Appendix D Ocean Current model
The current velocity expressed in the body-fixed frame is defined as
$\nu_{c}^{b}=\begin{bmatrix}u_{c},v_{c},w_{c}\end{bmatrix}^{\top}$. (61)
$\nu_{c}^{b}=\left(R_{b}^{n}\right)^{-1}{\nu}_{{c}}^{n}.$ (63)
$\dot{\nu}_{r}$ $=\dot{\nu}-\dot{\nu}_{c}$ (64)
$=M^{{\small-}1}\!\big(\tau\!+\!\tau_{LD}\!+\!\tau_{CF}{\small-}C(\nu_{r})\nu_{r}{\small-}D(\nu_{r})\nu_{r}{\small-}g(\eta)\big).$
Appendix E Hyperparameter Tuning Procedure
The DeePC HPs were tuned through a structured two-stage process to ensure robust performance across the full operating range of the AUV. In the first stage, broad parameter ranges were explored in an approximate search, where each HP was varied individually while the others were held constant. The resulting preliminary values formed the basis for the second stage, in which each HP was re-evaluated separately through ten independent simulations per candidate value—five under nominal conditions and five with external disturbances (ocean currents), due to the stochastic variability of the PRBS used for data collection. Different reference trajectories were applied across the repetitions to ensure robustness over the full operational range. Each trajectory contained an LP-smoothened step reference that was superimposed with a sinusoidal component, comparable to Fig. 5. The best-performing value was selected manually based on the RMSE of the overall tracking performance and qualitative assessment of control-input smoothness. Across all HPs, the control performance ”valley” was relatively flat and with a wide region of near-optimal values. Consequently, the final parameter set was chosen from within this region, prioritizing control smoothness over minor numerical differences in RMSE.

### v2

@@ 5 Data-Enabled Predictive Control
The DeePC framework, introduced in [7], relies on measured input–output trajectories that are sufficiently informative to capture the system dynamics, a property known as persistency of excitation [24]. According to Willems’ Fundamental Lemma, any trajectory of a controllable and observable LTI system can be expressed as a linear combination of previously recorded input–output data. The collected input and output sequences, $u_{\mathrm{data}}$ and $y_{\mathrm{data}}$, are arranged into block Hankel matrices $\mathscr{H}(u)$ and $\mathscr{H}(y)$ with $r$ block rows and $c$ columns, as defined in [7]. Alternatively, the data can be structured as a Page matrix $\mathscr{H}_{P}(u)$ as in  (57), which groups consecutive columns into separate pages [8].
@@
General persistency of excitation requires these data matrices to be of full row rank [7]. For noisy data, a quantitative characterization of this condition was introduced in [9], relating the degree of excitation to the singular value of the data matrix. For the DeePC formulation, the data matrices are partitioned into past and future components as in (58), where $(U_{p},Y_{p})$ encode past trajectories for initialization, and $(U_{f},Y_{f})$ define the future trajectories used for prediction.
@@
Here, $g\in\mathbb{R}^{T-T_{\mathrm{ini}}-N+1}$ is the decision vector that linearly combines past trajectories to reconstruct the system behavior, and $r\in\mathbb{R}^{p}$ denotes the reference trajectory. $Q\in\mathbb{R}^{p\times p}$ and $R\in\mathbb{R}^{m\times m}$ are positive definite weighting matrices. The regularization terms ${\|g\|_{\lambda_{g}}^{2}}$ and ${\|\sigma_{y}\|_{\lambda_{y}}^{2}}$, with $\lambda_{g},\lambda_{y}>0$, follow the approach of [17] to improve robustness against measurement noise and nonlinearities. Input and output constraints are directly incorporated through the box constraints and admissible sets $\mathcal{U}$ and $\mathcal{Y}$. Theoretical and experimental guarantees regarding robustness and stability are provided in [3, 4, 2].
@@ 6.2 DeePC Depth Control
While the data collection process for the inner loop is straightforward as for the heading control, collecting suitable data for the outer loop requires additional consideration, as its effective plant consists of the inner DeePC controller combined with the AUV dynamics. At first glance, this appears incompatible with the DeePC framework, as the inner DeePC involves an embedded optimization, breaking the deterministic LTI structure typically assumed by Willems’ Fundamental Lemma. To overcome this, a perfect control assumption, analogous to the unity-gain property in cascaded PID control [1], is introduced. This idealized condition assumes that the inner loop perfectly tracks its reference, i.e., $\theta=\theta_{d}$, thereby allowing the pitch angle $\theta$ to be used as the effective input and the measured heave position $z_{p}$ as the output for the outer-loop data sequences.
@@ 8 Results
This section presents the simulation results obtained for the proposed DeePC-based control framework and compares its performance with the classical PI/PID control architecture across various scenarios, including nominal operation, external disturbances, nonlinear dynamic conditions and 3-D path following. For the classical architecture in all following scenarios, the best-performing PI/PID gains were selected from the preinitialized MSS toolbox for depth control and the tuning proposed in [14] for heading control. The gain configuration of [13] was also tested but yielded inferior performance. The PI/PID gains are as follows:
[…]
@@ 8.1 General Control Performance
The PI/PID and DeePC-based controllers for heading and depth control are compared, with reference commands smoothed using the built-in LP filters of the MSS toolbox [12]. At $t=10\,\mathrm{s}$, a step command was applied toward a depth of $30\,\mathrm{m}$ and a heading of $60^{\circ}$. At $t=100\,\mathrm{s}$, a sinusoidal signal was superimposed on both reference trajectories to evaluate performance under more dynamic, albeit less realistic, reference variations. In both control approaches, the propeller speed was fixed to $n=1000\,\mathrm{rpm}$ during data collection and control execution. The specific DeePC hyperparameters (HP) are listed in Table 2.
@@ 8.3 Robustness Evaluation under Nonlinear Dynamics
8.5 Hyperparameter Tuning Procedure
The DeePC HPs were tuned through a structured two-stage process to ensure robust performance across the full operating range of the AUV. In the first stage, broad parameter ranges were explored in an approximate search, where each HP was varied individually while the others were held constant. The resulting preliminary values formed the basis for the second stage, in which each HP was re-evaluated separately through ten independent simulations per candidate value—five under nominal conditions and five with external disturbances (ocean currents), due to the stochastic variability of the PRBS used for data collection. Different reference trajectories were applied across the repetitions to ensure robustness over the full operational range. Each trajectory contained an LP-smoothened step reference that was superimposed with a sinusoidal component, comparable to Fig. 5. The best-performing value was selected manually based on the RMSE of the overall tracking performance and qualitative assessment of control-input smoothness. Across all HPs, the control performance ”valley” was relatively flat and with a wide region of near-optimal values. Consequently, the final parameter set was chosen from within this region, prioritizing control smoothness over minor numerical differences in RMSE. The selected HPs are summarized in Table 2. For the 3-D path-following case in Section 8.4, the $\theta$-DeePC controller uses slightly modified weights, specifically $Q=100$ and $R=100$.
@@ 9 Conclusion and Further Work
This work successfully demonstrated DeePC for AUV control in both heading and depth regulation in a high fidelity simulation environment. A cascaded DeePC architecture was developed for depth control, incorporating a loop-frequency separation concept that proved effective under disturbance and robustness tests. Furthermore, the PALOS algorithm was proposed and integrated with DeePC, enabling successful 3-D waypoint path following of the REMUS 100 AUV. The results confirm that the proposed PALOS-DeePC approach constitutes a robust, data-driven, and predictive control solution, achieving high performance without requiring an explicit dynamic model and the associated effort. DeePC consistently outperformed the classical PI/PID control architecture, and although more aggressive gain tuning could narrow the performance gap, the predictive nature of DeePC inherently offers superior anticipatory behavior compared to reactive controllers.
Future work will extend the approach towards more nonlinear operation across the full speed range with gain-scheduled DeePC formulations recently developed in [25], and advancing from waypoint-based path following to full trajectory tracking. Furthermore, experimental validation on a real AUV is planned to demonstrate the method’s practical applicability.